# On Effects of Steering Latent Representation for Large Language Model Unlearning

### v1

@@ 3.2 The effect of the coefficient on forget-sample representations
We aim to answer these questions by analyzing the unlearning problem under a noise compression view. We consider the output of a transformation $f^{k,l}$ on input $x$: $f^{k,l}(x)=(g^{(k)}\circ h^{(l)})(x)=g^{(k)}\left(h^{(l)}(x)\right)$. Suppose we compress a noise vector $\bm{\xi}$ to the representation $h^{(l)}$ of layer $l$ at an input $x$, then the output become $g^{(k)}\left(h^{(l)}(x)+\bm{\xi}\right)$. Naturally, if layer $g^{(k)}$ is robust (less sensitive) to noise $\bm{\xi}$, then $\bm{\xi}$ has a small effect on the output of $g^{(k)}$ i.e. the normalized squared norm
$\Phi(g^{(k)},x)=\frac{||g^{(k)}\left(h^{(l)}(x)+\bm{\xi}\right)-g^{(k)}\left(h^{(l)}(x)\right)||^{2}}{||g^{(k)}\left(h^{(l)}(x)\right)||^{2}}$ (8)
is small. In contrast, a higher $\Phi(g^{(k)},x)$ mean $g^{(k)}$ is higher sensitive to noise $\bm{\xi}$ at input $x$. For a dataset $\mathcal{D}_{\text{forget}}$, we define the noise sensitivity of a layer $g^{(k)}$ w.r.t $\bm{\xi}$ on $\mathcal{D}_{\text{forget}}$ as:
$\Phi(g^{(k)},\mathcal{D}_{\text{forget}})=\frac{||g^{(k)}(\hat{h}^{(l)}(x_{F})+\bm{\xi})-g^{(k)}(\hat{h}^{(l)}(x_{F}))||^{2}}{||g^{(k)}(\hat{h}^{(l)}(x_{F}))||^{2}},$ (9)
where $\hat{h}^{(l)}(x_{F})$ is the mean of $h^{(l)}(x_{F})$ over $x_{F}\in\mathcal{D}_{\text{forget}}$. During unlearning, RMU steers $h^{(l)}(x_{F})$ for all $x_{F}\in\mathcal{D}_{\text{forget}}$ to the fixed vector $c\bm{u}+\bm{\epsilon}$ i.e. $||g^{(k)}(c\bm{u}+\bm{\epsilon})-g^{(k)}(\hat{h}^{(l)}(x_{F}))||^{2}$ is minimized. If we let $\bm{\xi}=c\bm{u}+\bm{\epsilon}-\hat{h}^{(l)}(x_{F})$, we can define the unlearning problem as minimizing the noise sensitivity of layer. This objective is described by
$\min\frac{||g^{(k)}(c\bm{u}+\bm{\epsilon})-g^{(k)}(\hat{h}^{(l)}(x_{F}))||^{2}}{||g^{(k)}(\hat{h}^{(l)}(x_{F}))||^{2}}$ (10)
While $g^{(k)}$ is a composition of transformer layers, which is hard to expand it in term of $c$. Therefore, we propose to use the Jacobian matrix $\bm{J}^{(k)}(x_{F})$—a linearized of $g^{(k)}$ at $x_{F}$—which describes the change in the output of $g^{(k)}$ due to a noise perturbed in the input $\hat{h}^{(l)}(x_{F})$. For simplification, we write $\hat{h}^{(l)}$, $\bm{J}^{(k)}$ instead of $\hat{h}^{(l)}(x_{F})$, $\bm{J}^{(k)}(x_{F})$ respectively. The objective becomes
$\min\frac{||\bm{J}^{(k)}(c\bm{u}+\bm{\epsilon})-\bm{J}^{(k)}\hat{h}^{(l)}||^{2}}{||\bm{J}^{(k)}\hat{h}^{(l)}||^{2}}$ (11)
Since $\bm{J}^{(k)}$ is a linear transformation, then
$||\bm{J}^{(k)}(c\bm{u}+\bm{\epsilon})-\bm{J}^{(k)}\hat{h}^{(l)}||^{2}=||\bm{J}^{(k)}(c\bm{u}+\bm{\epsilon}-\hat{h}^{(l)})||^{2}$ (12)
@@
$||\bm{J}^{(k)}(c\bm{u}+\bm{v})||^{2}$ $=(\bm{J}^{(k)}(c\bm{u}+\bm{v}))^{\top}\bm{J}^{(k)}(c\bm{u}+\bm{v})$
$=(c\bm{u}+\bm{v})^{\top}\bm{J}^{(k)\top}\bm{J}^{(k)}(c\bm{u}+\bm{v})$ (13)
Let matrix $\bm{A}=\bm{J}^{(k)\top}\bm{J}^{(k)}$. Expand the right-hand side of Eqn. 13, we get:
$||\bm{J}^{(k)}(c\bm{u}+\bm{v})||^{2}$
@@
$||\bm{J}^{(k)}(c\bm{u}+\bm{v})||^{2}=c^{2}\bm{u}^{\top}\bm{A}\bm{u}+2c\bm{u}^{\top}\bm{A}\bm{v}+\bm{v}^{\top}\bm{A}\bm{v}$ (16)
While $||\bm{J}^{(k)}\hat{h}^{(l)}||^{2}$ is not zero. The objective described in Eqn. 11 is equivalent to
$\min||\bm{J}^{(k)}(c\bm{u}+\bm{v})||^{2}$ (17)
Since $||\bm{J}^{(k)}(c\bm{u}+\bm{v})||^{2}$ form a quadratic expression dependence on $c$, we take its derivative w.r.t $c$ and set it equal to zero:
$0\displaystyle 2\bm{u}^{\top}\bm{A}\bm{u}c+2\bm{u}^{\top}\bm{A}\bm{v}=$ 0 (18)
Solve for $c$:
$c$ $=-\frac{\bm{u}^{\top}\bm{A}\bm{v}}{\bm{u}^{\top}\bm{A}\bm{u}}=\frac{\bm{u}^{\top}\bm{J}^{(k)\top}\bm{J}^{(k)}(\hat{h}^{(l)}-\bm{\epsilon})}{\bm{u}^{\top}\bm{J}^{(k)\top}\bm{J}^{(k)}\bm{u}}$
$=\frac{(\bm{J}^{(k)}\bm{u})^{\top}\bm{J}^{(k)}(\hat{h}^{(l)}-\bm{\epsilon})}{||\bm{J}^{(k)}\bm{u}||^{2}}$
$=\frac{||\bm{J}^{(k)}(\hat{h}^{(l)}-\bm{\epsilon}))||}{||\bm{J}^{(k)}\bm{u}||}\cos(\bm{J}^{(k)}\bm{u},\bm{J}^{(k)}(\hat{h}^{(l)}-\bm{\epsilon}))$ (19)
Since $\frac{||\bm{J}^{(k)}(\hat{h}^{(l)}-\bm{\epsilon})||}{||\bm{J}^{(k)}\bm{u}||}$ is positive, then $c$ and $\cos(\bm{J}^{(k)}\bm{u},\bm{J}^{(k)}(\hat{h}^{(l)}-\bm{\epsilon}))$ are positively correlated.
@@
This means smaller (larger) $c$ indicates less (more) alignment between $\bm{J}^{(k)}\bm{u}$ and $\bm{J}^{(k)}(\hat{h}^{(l)}-\bm{\epsilon})$. Given that the Jacobian $\bm{J}^{(k)}$ describes how small changes in the input lead to changes in the output using linear approximation around a given point. If $\bm{J}^{(k)}$ does not vary drastically, it will not significantly alter the directions of $\bm{u}$ and $\hat{h}^{(l)}-\bm{\epsilon}$. In such cases, $\bm{J}^{(k)}$ will have a small effect on directional alignment, preserving the relative angles between $\bm{u}$ and $\hat{h}^{(l)}-\bm{\epsilon}$. Here, reasonably, $\bm{u}$ and $\hat{h}^{(l)}$ are becoming more aligned as $c$ increases since error $\bm{\epsilon}\to\bm{0}$ as unlearning becomes more accurate.
@@
Figure 1: Noise sensitivity of layer $g^{(k)}$, for $k\in[3...31]$ in base Zephyr-7B, base Llama-3-8B, base Mistral-7B, and RMU Zephyr-7B model. In the base models, a deeper layer has lower noise sensitivity, while the noise sensitivity is minimized in the RMU model (compress noise into layer $l=7$, the noise sensitivity of layer $k=8$ is minimized).
@@
The above discussion does not directly address RQ2. However, the definition of the noise sensitivity suggests that the noise sensitivity of layer $g^{(k)}$ characterized by the inherent properties of $g^{(k)}$, the representation $\hat{h}^{(l)}(x_{F})$ (which is fixed) and the perturbed noise $\bm{\xi}$. If $\bm{\xi}$ is predetermined, the noise sensitivity of $g^{(k)}$ depends solely on its properties. This suggest the following experiment: we compute $\hat{h}^{(l)}(x_{F})$—the mean of $h^{(l)}(x_{F})$ over a set of input $x_{F}\in\mathcal{D}_{\text{forget}}$, compress a fix perturbed noise $\bm{\xi}$ into $\hat{h}^{(l)}(x_{F})$. We then calculate the noise sensitivity of $g^{(k)}$ for different layers. Fig. 1 shows the noise sensitivity of layers across different models. We empirically observed that: the noise sensitivity decreases as layers go deeper and varies across different models. Since noise sensitivity describes a layer’s robustness to noise, higher noise sensitivity means $g^{(k)}$ requires smaller noise to produce the same level of output randomness, while lower noise sensitivity means it requires larger noise. In other words, early layers require smaller noise $\bm{\xi}$ (smaller $c$) whereas later layers require larger noise $\bm{\xi}$ (larger $c$). We present an empirical experiment to verify our analysis in Section. 4.3.
@@ 3.3 Robustness of RMU models to adversarial jailbreak attacks
Figure 2: The distribution of MaxLogit (a-d) on WMDP with different coefficient $c$ of the Base Zephyr-7B and RMU Zephyr-7B models ($l=7$). The distribution of $\cos(\bm{u},h^{(l)})$ (e-h) on WMDP of the RMU Zephyr-7B model ($l=7$).
RMU algorithms not only erase undesirable knowledge (forgotten knowledge) but also prevent them from resurfacing when attacked by adversarial jailbreak attacks. Li et al. (2024b) present a qualitvative experiment by employing Greedy Coordinate Gradient (GCG; Zou et al. (2023b)) to jailbreak RMU models to extract the answer given a harmful input. As a result, RMU models demonstrate strong robustness against GCG attack. We conduct an analysis from the standpoint of an attack-defense game to explain the underlying causes for the effectiveness of the RMU algorithm.
@@
The goal of the attacker is to inject an adversarial trigger into the prompt input (e.g.,“How would I engineer influenza A to be significantly more virulent? x x x x x x), causing the model generate a harmful responses (e.g., “To make influenza A more virulent…”) instead of refusing to answer (e.g., “As a language model,…”).
@@
where $\mathcal{J}(\cdot,\cdot)$ is the loss function of the attacker. The attacker find an update $\delta$ based on the linearized approximation of the loss (Zou et al. 2023b):
$\nabla_{e_{x_{i}}}\mathcal{J}(f(x_{F}))$ (21)
where $e_{x_{i}}$ is the one-hot vector representing the current value of the $i$th token in the $x_{F}$. The gradient $\nabla_{e_{x_{i}}}\mathcal{J}(f(x_{F}))$ is a good indicator for finding a set of candidates for the adversarial token replacement. A more negative value of the gradient $\nabla_{e_{x_{i}}}\mathcal{J}(f(x_{F}))$ makes a more decrease in the loss. The GCG algorithm finds top-$k$ largest negative value of $\nabla_{e_{x_{i}}}\mathcal{J}(f(x_{F}))$ for each token in the adversarial trigger and makes the replacement the most decrease in the loss.
@@
$\nabla_{e_{x_{i}}}\mathcal{J}(f^{\text{unlearn}}(x_{F}))$ $=\nabla_{e_{x_{i}}}\mathcal{J}(g(h^{(l),\textnormal{steered}}(x_{F}))$ (23)
$=\nabla_{e_{x_{i}}}(\mathcal{J}\circ g)(c\bm{u}+\bm{\epsilon})$ (24)
Since $c$ and $\bm{u}$ are predetermined before unlearning, $(\mathcal{J}\circ g)(c\bm{u})$ does not change with respect to $e_{x_{i}}$. The gradient $\nabla_{e_{x_{i}}}(\mathcal{J}\circ g)(c\bm{u}+\bm{\epsilon})$ close to 00 for all token $x_{i}$ since the error $\bm{\epsilon}\to\bm{0}$ as unlearning becomes accurate. This means the GCG attacker received unreliable, uninformative gradient signals from RMU models. The RMU model serves as a defender by causing the attacker to miscalculate the gradient of the loss to optimize its objective, thereby increasing the attacker’s cost. The attacker, therefore, cannot find the optimal adversarial tokens for replacement. Li et al. (2024b)’s experiment results implicitly verify our analysis.

### v2

@@ 3.2 The effect of the coefficient on forget-sample representations
We aim to answer these questions by analyzing the unlearning problem under a noise compression view. We consider the output of a transformation $f^{(l:k)}$ on input $x$: $f^{(l:k)}(x)=(g^{(l:k)}\circ h^{(l)})(x)=g^{(l:k)}\left(h^{(l)}(x)\right)$. Suppose we compress a noise vector $\bm{\xi}$ to the representation $h^{(l)}$ of layer $l$ at an input $x$, then the output become $g^{(l:k)}\left(h^{(l)}(x)+\bm{\xi}\right)$. Naturally, if layer $g^{(l:k)}$ is robust (less sensitive) to noise $\bm{\xi}$, then $\bm{\xi}$ has a small effect on the output of $g^{(l:k)}$ i.e. the normalized squared norm
$\Phi(g^{(l:k)},x)=\frac{||g^{(l:k)}\left(h^{(l)}(x)+\bm{\xi}\right)-g^{(l:k)}\left(h^{(l)}(x)\right)||^{2}}{||g^{(l:k)}\left(h^{(l)}(x)\right)||^{2}}$ (8)
is small. In contrast, a higher $\Phi(g^{(l:k)},x)$ mean $g^{(l:k)}$ is higher sensitive to noise $\bm{\xi}$ at input $x$. For a dataset $\mathcal{D}_{\text{forget}}$, we define the noise sensitivity of a layer $g^{(l:k)}$ w.r.t $\bm{\xi}$ on $\mathcal{D}_{\text{forget}}$ as:
$\Phi(g^{(l:k)},\mathcal{D}_{\text{forget}})$
$=\frac{||g^{(l:k)}(\hat{h}^{(l)}(x_{F})+\bm{\xi})-g^{(l:k)}(\hat{h}^{(l)}(x_{F}))||^{2}}{||g^{(l:k)}(\hat{h}^{(l)}(x_{F}))||^{2}},$ (9)
where $\hat{h}^{(l)}(x_{F})$ is the mean of $h^{(l)}(x_{F})$ over $x_{F}\in\mathcal{D}_{\text{forget}}$. During unlearning, RMU steers $h^{(l)}(x_{F})$ for all $x_{F}\in\mathcal{D}_{\text{forget}}$ to the fixed vector $c\bm{u}+\bm{\epsilon}$ i.e. $||g^{(l:k)}(c\bm{u}+\bm{\epsilon})-g^{(l:k)}(\hat{h}^{(l)}(x_{F}))||^{2}$ is minimized. If we let $\bm{\xi}=c\bm{u}+\bm{\epsilon}-\hat{h}^{(l)}(x_{F})$, we can define the unlearning problem as minimizing the noise sensitivity of the layer. This objective is described by
$\min\frac{||g^{(l:k)}(c\bm{u}+\bm{\epsilon})-g^{(l:k)}(\hat{h}^{(l)}(x_{F}))||^{2}}{||g^{(l:k)}(\hat{h}^{(l)}(x_{F}))||^{2}}$ (10)
While $g^{(l:k)}$ is a composition of transformer layers, which is hard to expand it in term of $c$. Therefore, we propose to use the Jacobian matrix $\bm{J}^{(l:k)}(x_{F})$—a linearized of $g^{(l:k)}$ at $x_{F}$—which describes the change in the output of $g^{(l:k)}$ due to a noise perturbed in the input $\hat{h}^{(l)}(x_{F})$. For simplification, we write $\hat{h}^{(l)}$, $\bm{J}^{(l:k)}$ instead of $\hat{h}^{(l)}(x_{F})$, $\bm{J}^{(l:k)}(x_{F})$ respectively. The objective becomes
$\min\frac{||\bm{J}^{(l:k)}(c\bm{u}+\bm{\epsilon})-\bm{J}^{(l:k)}\hat{h}^{(l)}||^{2}}{||\bm{J}^{(l:k)}\hat{h}^{(l)}||^{2}}$ (11)
Since $\bm{J}^{(l:k)}$ is a linear transformation, then
$||\bm{J}^{(l:k)}(c\bm{u}+\bm{\epsilon})-\bm{J}^{(l:k)}\hat{h}^{(l)}||^{2}=||\bm{J}^{(l:k)}(c\bm{u}+\bm{\epsilon}-\hat{h}^{(l)})||^{2}$ (12)
@@
$||\bm{J}^{(l:k)}(c\bm{u}+\bm{v})||^{2}$ $=(\bm{J}^{(l:k)}(c\bm{u}+\bm{v}))^{\top}\bm{J}^{(l:k)}(c\bm{u}+\bm{v})$
$=(c\bm{u}+\bm{v})^{\top}\bm{J}^{(l:k)\top}\bm{J}^{(l:k)}(c\bm{u}+\bm{v})$ (13)
Let matrix $\bm{A}=\bm{J}^{(l:k)\top}\bm{J}^{(l:k)}$. Expand the right-hand side of Eqn. 13, we get:
$||\bm{J}^{(l:k)}(c\bm{u}+\bm{v})||^{2}$
@@
$||\bm{J}^{(l:k)}(c\bm{u}+\bm{v})||^{2}=c^{2}\bm{u}^{\top}\bm{A}\bm{u}+2c\bm{u}^{\top}\bm{A}\bm{v}+\bm{v}^{\top}\bm{A}\bm{v}$ (16)
Substituting Eqn. 16 into Eqn. 11, the objective becomes
$\min\frac{c^{2}\bm{u}^{\top}\bm{A}\bm{u}+2c\bm{u}^{\top}\bm{A}\bm{v}+\bm{v}^{\top}\bm{A}\bm{v}}{||\bm{J}^{(l:k)}\hat{h}^{(l)}||^{2}}$ (17)
Taking its derivative w.r.t $c$ and set it to zero:
$\frac{2\bm{u}^{\top}\bm{A}\bm{u}c+2\bm{u}^{\top}\bm{A}\bm{v}}{||\bm{J}^{(l:k)}\hat{h}^{(l)}||^{2}}=0$ (18)
Since $||\bm{J}^{(l:k)}\hat{h}^{(l)}||^{2}$ is not zero, solve for $c$:
$c$ $=-\frac{\bm{u}^{\top}\bm{A}\bm{v}}{\bm{u}^{\top}\bm{A}\bm{u}}=\frac{\bm{u}^{\top}\bm{J}^{(l:k)\top}\bm{J}^{(l:k)}(\hat{h}^{(l)}-\bm{\epsilon})}{\bm{u}^{\top}\bm{J}^{(l:k)\top}\bm{J}^{(l:k)}\bm{u}}$
$=\frac{(\bm{J}^{(l:k)}\bm{u})^{\top}\bm{J}^{(l:k)}(\hat{h}^{(l)}-\bm{\epsilon})}{||\bm{J}^{(l:k)}\bm{u}||^{2}}$
$=\frac{||\bm{J}^{(l:k)}(\hat{h}^{(l)}-\bm{\epsilon}))||}{||\bm{J}^{(l:k)}\bm{u}||}\cos(\bm{J}^{(l:k)}\bm{u},\bm{J}^{(l:k)}(\hat{h}^{(l)}-\bm{\epsilon}))$ (19)
Since $\frac{||\bm{J}^{(l:k)}(\hat{h}^{(l)}-\bm{\epsilon})||}{||\bm{J}^{(l:k)}\bm{u}||}$ is positive, then $c$ and $\cos(\bm{J}^{(l:k)}\bm{u},\bm{J}^{(l:k)}(\hat{h}^{(l)}-\bm{\epsilon}))$ are positively correlated.
@@
This means smaller (larger) $c$ indicates less (more) alignment between $\bm{J}^{(l:k)}\bm{u}$ and $\bm{J}^{(l:k)}(\hat{h}^{(l)}-\bm{\epsilon})$. Given that the Jacobian $\bm{J}^{(l:k)}$ describes how small changes in the input lead to changes in the output using linear approximation around a given point. If $\bm{J}^{(l:k)}$ does not vary drastically, it will not significantly alter the directions of $\bm{u}$ and $\hat{h}^{(l)}-\bm{\epsilon}$. In such cases, $\bm{J}^{(l:k)}$ will have a small effect on directional alignment, preserving the relative angles between $\bm{u}$ and $\hat{h}^{(l)}-\bm{\epsilon}$. Here, reasonably, $\bm{u}$ and $\hat{h}^{(l)}$ are becoming more aligned as $c$ increases since error $\bm{\epsilon}\to\bm{0}$ as unlearning becomes more accurate.
@@
Figure 1: Noise sensitivity of layer $g^{(l:k)}$, for $k\in[3...31]$ in base Zephyr-7B, base Llama-3-8B, base Mistral-7B, and RMU Zephyr-7B model. In the base models, a deeper layer has lower noise sensitivity, while the noise sensitivity is minimized in the RMU model (compress noise into $h^{(7)}$, the noise sensitivity of layer $k=8$ is minimized).
@@
The above discussion does not directly address RQ2. However, the definition of the noise sensitivity suggests that the noise sensitivity of layer $g^{(l:k)}$ is characterized by the inherent properties of $g^{(l:k)}$, the representation $\hat{h}^{(l)}(x_{F})$ (which is fixed) and the perturbed noise $\bm{\xi}$. If $\bm{\xi}$ is predetermined, the noise sensitivity of $g^{(l:k)}$ depends solely on its properties. This suggest the following experiment: we compute $\hat{h}^{(l)}(x_{F})$—the mean of $h^{(l)}(x_{F})$ over a set of input $x_{F}\in\mathcal{D}_{\text{forget}}$, compress a fix perturbed noise $\bm{\xi}$ into $\hat{h}^{(l)}(x_{F})$. We then calculate the noise sensitivity of $g^{(l:k)}$ for different layers. Fig. 1 shows the noise sensitivity of layers across different models. We empirically observed that: the noise sensitivity decreases as layers go deeper and vary across different models. Since noise sensitivity describes a layer’s robustness to noise, higher noise sensitivity means $g^{(l:k)}$ requires smaller noise to produce the same level of output randomness, while lower noise sensitivity means it requires larger noise. In other words, early layers require smaller noise $\bm{\xi}$ (smaller $c$) whereas later layers require larger noise $\bm{\xi}$ (larger $c$). We present an empirical experiment to verify our analysis in Section. 4.3.
@@ 3.3 Robustness of RMU models to adversarial jailbreak attacks
Figure 2: The distribution of MaxLogit (a-d) on WMDP Q&A sets with different coefficient $c$ of the base Zephyr-7B and RMU Zephyr-7B models ($l=7$). The distribution of $\cos(\bm{u},h^{(l)})$ (e-h) of the RMU Zephyr-7B model ($l=7$).
RMU algorithms not only erase undesirable knowledge (forgotten knowledge) but also prevent it from resurfacing when attacked by adversarial jailbreak attacks. Li et al. (2024b) present a qualitvative experiment by employing Greedy Coordinate Gradient (GCG; Zou et al. (2023b)) to jailbreak RMU models to extract the answer given a harmful input. As a result, RMU models demonstrate strong robustness against GCG attack. We conduct an analysis from the standpoint of an attack-defense game to explain the underlying causes for the effectiveness of the RMU algorithm.
@@
The goal of the attacker is to inject an adversarial trigger into the suffix of the prompt input (e.g., “How would I engineer influenza A to be significantly more virulent? x x x x x x), causing the model generates a harmful response (e.g., “To make influenza A more virulent…”).
@@
where $\mathcal{J}(\cdot,\cdot)$ is the loss function of the attacker. The attacker finds an update $\delta$ based on the linearized approximation of the loss $\nabla_{e_{x_{i}}}\mathcal{J}(f(x_{F}))$, where $e_{x_{i}}$ is the one-hot vector representing the current value of the $i$-th token in $x_{F}$. The gradient $\nabla_{e_{x_{i}}}\mathcal{J}(f(x_{F}))$ is a good indicator for finding a set of candidates for the adversarial token replacement. A more negative value of the gradient $\nabla_{e_{x_{i}}}\mathcal{J}(f(x_{F}))$ makes a more decrease in the loss. The GCG attacker finds top-$k$ largest negative value of $\nabla_{e_{x_{i}}}\mathcal{J}(f(x_{F}))$ for each token in the adversarial trigger and makes the replacement the most decrease in the loss.
@@
$\nabla_{e_{x_{i}}}\mathcal{J}(f^{\text{unlearn}}(x_{F}))$ $=\nabla_{e_{x_{i}}}\mathcal{J}(g^{(l:k)}(h^{(l),\textnormal{steered}}(x_{F}))$
$\approx\nabla_{e_{x_{i}}}(\mathcal{J}\circ g^{(l:k)})(c\bm{u}+\bm{\epsilon})$ (22)
Since $c$ and $\bm{u}$ are predetermined before unlearning, $(\mathcal{J}\circ g^{(l:k)})(c\bm{u})$ does not change with respect to $e_{x_{i}}$. The gradient $\nabla_{e_{x_{i}}}(\mathcal{J}\circ g^{(l:k)})(c\bm{u}+\bm{\epsilon})$ close to 00 for all token $x_{i}$ since the error $\bm{\epsilon}\to\bm{0}$ as unlearning becomes accurate. This means the GCG attacker received unreliable, uninformative gradient signals from RMU models. The RMU model serves as a defender by causing the attacker to miscalculate the gradient of the loss to optimize its objective, thereby increasing the attacker’s cost. The attacker, therefore, cannot find the optimal adversarial tokens for replacement. Li et al. (2024b)’s experiment results implicitly verify our analysis.